# Late-blooming magnetars: awakening as long period transients after a dormant cooling epoch

### v2

@@ III.2 X-rays and interpulses
Our CT energy distributions imply dim outbursts: if $\sim 1\%$ of the failure energy converts to observable X-rays (Blaes et al., 1989; Bransgrove, Beloborodov & Levin, 2020), most bursts would fall below detection thresholds. This may explain why only DA J1832 has shown coincident X-ray and radio activity, reflecting observational limitations rather than an intrinsic rarity of outbursts. Moreover, many LPTs were identified only in archival data, and thus X-ray follow-up some years after radio activation is inappropriate to assess the scenario proposed here. In contrast, high-cadence monitoring of the relatively active (Hurley-Walker et al., 2023; Men et al., 2025) LPT GPM J1839–10 ($D\sim 25\%$) would be especially useful to test the ‘late-bloomer’ picture.
@@ IV Enhanced spindown
Neutron stars gradually spin down due to electromagnetic braking torques. For a star with a (time-dependent) polar-dipole strength $B_{p}$, inclined relative to the rotation axis by an angle $\alpha$, the period evolves according to [^1] Note that we ignore general-relativistic corrections, which accelerate spindown for a given $B_{p}$, to provide a conservative scenario (see, e.g., Stefanou, Suvorov & Pons, 2025). (Spitkovsky, 2006)
@@
for moment of inertia $I_{0}$. However, if the external field anchored to the crust is twisted by failures, the resulting increase in magnetic pressure will open up more lines that carry away angular momentum (Thompson, Lyutikov & Kulkarni, 2002; Beloborodov, 2009). This will occur in stages, with a more ‘glitch-like’ jump in $P$ caused by the abrupt release followed by a gradual deceleration as the twist dissipates. Evidence for a timing glitch in the LPT CHIME J0630 supports this picture (Dong et al., 2025); sadly, the source was not monitored in X-rays at the time to compare with the model.
@@
where $E_{\rm rot}=2\pi^{2}I_{0}/P^{2}$. An estimate for the efficiency can be made by first comparing theoretical expectations for how much of the failure energy escapes as X-rays, thought to be of order $\sim 1\%$ (Blaes et al., 1989; Bransgrove, Beloborodov & Levin, 2020). Observations of spindown losses (Borghese et al., 2021; Lower et al., 2023) beyond the baseline model (7) following outbursts from radio-loud magnetars with pulse profiles most similar to that of LPTs indicate that $\Delta E_{\rm rot}\sim-0.1E_{\rm burst}$. The overall efficiency is the product of these two fractions, and so we consider $\bar{\zeta}=10^{-3}$ as representative (see Appendix C for further details on this choice).
@@
While CC models dominated by small-scales and with stronger fields can better explain X-ray observations of Galactic magnetars (Fig. 2), they fare worse in terms of explaining LPTs. This is because, even though $\mathcal{R}_{m}\propto B$ is itself independent of the system length-scale, both the Hall and Ohmic times scale as $L^{2}$ which is smaller for multipoles (equation 2). This prevents any CC model from reaching LPT-like periods even if $P\sim 10$ s earlier in life due to the stronger field (Pons, Viganò & Rea, 2013). In CT simulations, failures only start taking place en masse when $P\sim 10$ s and $\mathcal{R}_{m}\gtrsim 100$. Around $\gtrsim 1$ Myr, the cumulative failure energy becomes comparable to the rotational energy for $\bar{\zeta}\approx 10^{-3}$, indicating that injected twists could effectively drain the entire rotational kinetic energy reservoir.
@@
The amplitude of this enhancement is parameterised by the dimensionless efficiency factor, $\zeta$, defined in Eq. (8). The normalisation of this factor is constrained by observations: we require that the resulting spindown rate satisfies $\dot{P}<10^{-9}\mathrm{\,ss^{-1}}$, consistent with the observational upper limit inferred for DA J1832 (Wang et al., 2025). Following the rise phase, the enhanced dipole moment is assumed to decay linearly back to its pre-event value, in rough agreement with the observed relaxation of magnetospheric twists following magnetar outbursts (Turolla, Zane & Watts, 2015). The decay timescale is not chosen arbitrarily, but is fixed self-consistently by enforcing Eq. (8) as an integral constraint on the total rotational energy extracted during each event (i.e., that the time-integrated energies match).
@@ V Discussion and future prospects
We argue in this paper that the defining characteristics of isolated LPTs can be explained with a magnetar model provided that electrical currents circulate through the core (a lá CT). This condition alters the star’s trajectory in the $T$-$B$ phase space, as illustrated in Fig. 1 depicting the magnetic Reynolds number. Our simulations show that $\mathcal{R}_{m}\gtrsim 10^{2}$ at $\gtrsim 0.1$ Myr ages, whereupon the star is cold ($T_{s}\lesssim 10^{6}$ K) and too dim to be observable in X-rays (Fig. 2). At this age, it enters into a delayed Hall-dominated epoch such that it effectively reawakens (Pons & Geppert, 2007; Viganò et al., 2013) as what we have called a ‘late-blooming magnetar’ whereupon crustal failures (Fig. 3) start seeding both radio and X-ray activity. Critically, rotational energy is already low ($\sim 10^{44}$ erg; Fig. 4) at this stage meaning that, if even a fraction ($\bar{\zeta}\sim 10^{-3}$; Fig. 5) of the failure energy goes into spindown, the star could reach hour-long periods. The radio duty cycle is tied to the failure rate in this picture; a prediction of the model is thus that of a positive correlation between periods and duty cycles in LPTs. This is supported by the (small sample-size) data as the lowest-$P$ source (CHIME J0630) also has the lowest duty cycle ($D\sim 0.6\%$). GPM J1839–10 is an anomaly however, and we require that its pole is face-on so that the observed duty cycle is about half of the physical one ($D\approx D^{\rm p}/2$) or that it is in a binary (see Suvorov, Dehman & Pons, 2025, for a discussion).
@@
We find that failures cluster around the poles if the birth field is well-ordered, naturally explaining the latitude of the hotspot of DA J1832 (Wang et al., 2025) and interpulses from ASKAP J1839 (Lee et al., 2025). We note, however, that this polar clustering is also influenced by the axisymmetric nature of the simulations, which favours the development of coherent large-scale structures, as reported in earlier 2D studies (Perna & Pons, 2011). Fully 3D models may broaden the latitude distribution of failure sites, although we expect polar regions to remain favoured when the large-scale field geometry is predominantly dipolar.
@@
Low multipolarity also favours a fossil origin over a turbulent dynamo for the birth field (Barrère et al., 2022). For example, recent 3D simulations of core-collapse by Varma, Müller & Schneider (2023) for a static, magnetised $17M_{\odot}$-star found that $0.4$ s after bounce the dipole field ($\approx 3\times 10^{14}$ G) could make up $\approx 40\%$ of the RMS intensity ($\approx 7\times 10^{14}$ G) at densities surrounding $10^{11}\text{ g cm}^{-3}$. Promising progenitors are stars like $\xi^{1}$ Canis Majoris, a highly-magnetised $\sim 14M_{\odot}$ variable revolving once every >30>30 years (Shultz et al., 2017). If the star continues to slow at its current rate it would be static in <<4 Myr, whereupon collapse a slow magnetar may form by magnetic-flux and angular-momentum conservation. This scenario fits best with the RCW 103 supernova remnant 1E 161348–5055, which exhibits a variable X-ray luminosity spanning $\sim 10^{33}$ to $\sim 10^{35}$ erg/s indicative of ongoing heating despite showing a period of many hours (De Luca et al., 2006). This object is difficult to explain with a CT model, but failures in a CC model could drain the entire reservoir after a few kyr if it were born with $E_{\rm rot}\lesssim 10^{45}$ erg (Fig. 3).
@@
Mass-loaded winds could also lead to exponential increases in the spin period in a manner strongly resembling our Fig. 5 (see Harding, Contopoulos & Kazanas, 1999; Beniamini, Wadiasingh & Metzger, 2020; Beniamini et al., 2023). While previous models of flare-induced spindown are phenomenologically similar to that proposed here, we emphasise that our predictions (i) arise self-consistently from magnetothermal outputs, (ii) do not rely on global events, and (iii) only work in CT configurations for typical birth periods. The second point is worth bearing in mind since, although extensive torque variability was observed in the lead-up and aftermath to the event, minimal changes were seen in $\dot{P}$ surrounding the giant flare from SGR 1806–20 (Younes et al., 2017). One way in which these scenarios could be confronted in future is through metal abundances: mass-loaded winds should invite $r$-process nucleosynthesis and hence enrich LPT nurseries (Patel et al., 2025).
@@
Another possibility proposed by Lander et al. (2026) for producing isolated LPTs is that of a delayed Meissner expulsion. This could resurrect a slow, CC magnetar by pushing flux into the crust. This scenario would, in essence, be represented by the left panel of Fig. 4 but with a lower ‘initial’ $E_{\rm rot}$, so that when crustal failures begin to sap energy the star could reach $\bar{\zeta}E^{\rm acc}_{\rm fail}>E_{\rm rot}$. However, forcing huge fluxes into the crust should generate heat: this fits with 1E 1613 but is difficult to reconcile with the coldest LPTs. One may also anticipate a skew towards larger burst energies in this case (top panel of Fig. 3), which is tension with the low energy of the DA J1832 event and lack of visible outbursts from other LPTs. If a bright flare from a LPT is detected in future, such a scenario may be favoured since the maximum failure energy of CT models is of order $\sim 10^{41}$ erg (bottom panel of Fig. 3). Hall waves of high amplitude may also be launched into the crust as flux is thrust there (Bransgrove, Levin & Beloborodov, 2025), through a Meissner expulsion or otherwise. Notably, waves of the low-frequency, whistler variety may enduringly propagate in latitudinal directions such that the ambient poloidal field oscillates violently, instigating intermittent heating, anomalous torques, avalanche phenomena, and multiband transients. Ab initio modelling in this direction could pave the way for population synthesis studies of canonical vs. ‘late-blooming’ vs. ‘reborn’ magnetars.
@@ Acknowledgements
AGS thanks Adam Dong, Zorawar Wadiasingh, Yuri Levin, and Andrei Beloborodov for valuable discussions. We thank Alex Cooper for insights on pulsar activity driven by plastic flow. AGS acknowledges funding from the European Union’s Horizon MSCA-2022 research and innovation programme “EinsteinWaves” under grant agreement No. 101131233, the Deutsche Forschungsgemeinschaft individual research grant 570901071, and the High Performance and Cloud Computing Group at the Zentrum für Datenverarbeitung of the University of Tübingen (Project “AnqaGW”). CD is supported by a Juan de la Cierva fellowship JDC2023‐052227‐I, funded by MINISTERIO DE CIENCIA, INNOVACIÓN Y UNIVERSIDADES, AGENCIA ESTATAL DE INVESTIGACIÓN, and UNIÓN EUROPEA (FSE+). We gratefully acknowledge the support provided by the Conselleria d’Educació, Cultura, Universitats i Ocupació de la Generalitat Valenciana through Prometeo Project CIPROM/2022/13, and from the AEI grant PID2021-127495NB-I00 funded by MCIN/AEI/10.13039/501100011033.
@@ Appendix A Numerical details on magnetothermal scheme
We incorporate superfluid and superconducting gaps for neutrons and protons, respectively, as they play a crucial role in cooling timescales by influencing the heat capacity and neutrino emissivity. While our model includes these effects, we do not alter the equations of motion to incorporate additional effects. For a comprehensive formulation of dissipative magnetohydrodynamic equations for finite-temperature superfluid and superconducting charged relativistic mixtures that account for particle diffusion, we refer to Dommes & Gusakov (2021). Thermal and electrical conductivities are computed using realistic data from the IOFFE repository [^2] http://www.ioffe.ru/astro/conduct/. For a complete discussion of all these microphysical ingredients, beyond that which we can adequately explore here, we refer the reader to Potekhin, Pons & Page (2015).
To describe the hydrostatics of the background state, we solve the Tolman-Oppenheimer-Volkov equations with the BSk24 equation of state (EOS), not accounting for (small) thermal corrections (Pearson et al., 2018). This EOS includes consistent crust and core regions and matches well with observational constraints from gravitational waves from binary-mergers, maximum-mass observations of pulsars, fast-cooling constraints, phase-coherent X-ray spectra mapped by the Neutron Star Interior Composition Explorer Mission (NICER), and other channels (Suvorov, Kuan & Kokkotas, 2024). At the magnetic field strengths and temperatures considered here, such a cold and unmagnetised EOS is an excellent approximation. A spherically-symmetric star is assumed, ignoring rotational oblateness and/or magnetic deformations to the mass-density. The above factors could, in principle, influence the evolution (at least at early times) and introduce complicated anisotropies into the crustal conductivity profiles, but such effects are expected to be small.

### v3

@@ III.2 X-rays and interpulses
Our CT energy distributions imply dim outbursts: if $\sim 1\%$ of the failure energy converts to observable X-rays (Blaes et al., 1989; Bransgrove et al., 2020), most bursts would fall below detection thresholds. This may explain why only DA J1832 has shown coincident X-ray and radio activity, reflecting observational limitations rather than an intrinsic rarity of outbursts. Moreover, many LPTs were identified only in archival data, and thus X-ray follow-up some years after radio activation is inappropriate to assess the scenario proposed here. In contrast, high-cadence monitoring of the relatively active (Hurley-Walker et al., 2023; Men et al., 2025) LPT GPM J1839–10 ($D\sim 25\%$) would be especially useful to test the ‘late-bloomer’ picture.
@@ IV Enhanced spindown
Neutron stars gradually spin down due to electromagnetic braking torques. For a star with a (time-dependent) polar-dipole strength $B_{p}$, inclined relative to the rotation axis by an angle $\alpha$, the period evolves according to [^2] Note that we ignore general-relativistic corrections, which accelerate spindown for a given $B_{p}$, to provide a conservative scenario (see, e.g., Stefanou et al., 2025). (Spitkovsky, 2006)
@@
for moment of inertia $I_{0}$. However, if the external field anchored to the crust is twisted by failures, the resulting increase in magnetic pressure will open up more lines that carry away angular momentum (Thompson et al., 2002; Beloborodov, 2009). This will occur in stages, with a more ‘glitch-like’ jump in $P$ caused by the abrupt release followed by a gradual deceleration as the twist dissipates. Evidence for a timing glitch in the LPT CHIME J0630 supports this picture (Dong et al., 2025); sadly, the source was not monitored in X-rays at the time to compare with the model.
@@
where $E_{\rm rot}=2\pi^{2}I_{0}/P^{2}$. An estimate for the efficiency can be made by first comparing theoretical expectations for how much of the failure energy escapes as X-rays, thought to be of order $\sim 1\%$ (Blaes et al., 1989; Bransgrove et al., 2020). Observations of spindown losses (Borghese et al., 2021; Lower et al., 2023) beyond the baseline model (7) following outbursts from radio-loud magnetars with pulse profiles most similar to that of LPTs indicate that $\Delta E_{\rm rot}\sim-0.1E_{\rm burst}$. The overall efficiency is the product of these two fractions, and so we consider $\bar{\zeta}=10^{-3}$ as representative (see Appendix C for further details on this choice).
@@
While CC models dominated by small-scales and with stronger fields can better explain X-ray observations of Galactic magnetars (Fig. 2), they fare worse in terms of explaining LPTs. This is because, even though $\mathcal{R}_{m}\propto B$ is itself independent of the system length-scale, both the Hall and Ohmic times scale as $L^{2}$ which is smaller for multipoles (equation 2). This prevents any CC model from reaching LPT-like periods even if $P\sim 10$ s earlier in life due to the stronger field (Pons et al., 2013). In CT simulations, failures only start taking place en masse when $P\sim 10$ s and $\mathcal{R}_{m}\gtrsim 100$. Around $\gtrsim 1$ Myr, the cumulative failure energy becomes comparable to the rotational energy for $\bar{\zeta}\approx 10^{-3}$, indicating that injected twists could effectively drain the entire rotational kinetic energy reservoir.
@@
The amplitude of this enhancement is parameterised by the dimensionless efficiency factor, $\zeta$, defined in Eq. (8). The normalisation of this factor is constrained by observations: we require that the resulting spindown rate satisfies $\dot{P}<10^{-9}\mathrm{\,ss^{-1}}$, consistent with the observational upper limit inferred for DA J1832 (Wang et al., 2025). Following the rise phase, the enhanced dipole moment is assumed to decay linearly back to its pre-event value, in rough agreement with the observed relaxation of magnetospheric twists following magnetar outbursts (Turolla et al., 2015). The decay timescale is not chosen arbitrarily, but is fixed self-consistently by enforcing Eq. (8) as an integral constraint on the total rotational energy extracted during each event (i.e., that the time-integrated energies match).
@@ V Discussion and future prospects
We argue in this paper that the defining characteristics of isolated LPTs can be explained with a magnetar model provided that electrical currents circulate through the core (a lá CT). This condition alters the star’s trajectory in the $T$-$B$ phase space, as illustrated in Fig. 1 depicting the magnetic Reynolds number. Our simulations show that $\mathcal{R}_{m}\gtrsim 10^{2}$ at $\gtrsim 0.1$ Myr ages, whereupon the star is cold ($T_{s}\lesssim 10^{6}$ K) and too dim to be observable in X-rays (Fig. 2). At this age, it enters into a delayed Hall-dominated epoch such that it effectively reawakens (Pons and Geppert, 2007; Viganò et al., 2013) as what we have called a ‘late-blooming magnetar’ whereupon crustal failures (Fig. 3) start seeding both radio and X-ray activity. Critically, rotational energy is already low ($\sim 10^{44}$ erg; Fig. 4) at this stage meaning that, if even a fraction ($\bar{\zeta}\sim 10^{-3}$; Fig. 5) of the failure energy goes into spindown, the star could reach hour-long periods. The radio duty cycle is tied to the failure rate in this picture; a prediction of the model is thus that of a positive correlation between periods and duty cycles in LPTs. This is supported by the (small sample-size) data as the lowest-$P$ source (CHIME J0630) also has the lowest duty cycle ($D\sim 0.6\%$). GPM J1839–10 is an anomaly however, and we require that its pole is face-on so that the observed duty cycle is about half of the physical one ($D\approx D^{\rm p}/2$) or that it is in a binary (see Suvorov et al., 2025, for a discussion).
@@
We find that failures cluster around the poles if the birth field is well-ordered, naturally explaining the latitude of the hotspot of DA J1832 (Wang et al., 2025) and interpulses from ASKAP J1839 (Lee et al., 2025). We note, however, that this polar clustering is also influenced by the axisymmetric nature of the simulations, which favours the development of coherent large-scale structures, as reported in earlier 2D studies (Perna and Pons, 2011). Fully 3D models may broaden the latitude distribution of failure sites, although we expect polar regions to remain favoured when the large-scale field geometry is predominantly dipolar.
@@
Low multipolarity also favours a fossil origin over a turbulent dynamo for the birth field (Barrère et al., 2022). For example, recent 3D simulations of core-collapse by Varma et al. (2023) for a static, magnetised $17M_{\odot}$-star found that $0.4$ s after bounce the dipole field ($\approx 3\times 10^{14}$ G) could make up $\approx 40\%$ of the RMS intensity ($\approx 7\times 10^{14}$ G) at densities surrounding $10^{11}\text{ g cm}^{-3}$. Promising progenitors are stars like $\xi^{1}$ Canis Majoris, a highly-magnetised $\sim 14M_{\odot}$ variable revolving once every >30>30 years (Shultz et al., 2017). If the star continues to slow at its current rate it would be static in <<4 Myr, whereupon collapse a slow magnetar may form by magnetic-flux and angular-momentum conservation. This scenario fits best with the RCW 103 supernova remnant 1E 161348–5055, which exhibits a variable X-ray luminosity spanning $\sim 10^{33}$ to $\sim 10^{35}$ erg/s indicative of ongoing heating despite showing a period of many hours (De Luca et al., 2006). This object is difficult to explain with a CT model, but failures in a CC model could drain the entire reservoir after a few kyr if it were born with $E_{\rm rot}\lesssim 10^{45}$ erg (Fig. 3).
@@
Mass-loaded winds could also lead to exponential increases in the spin period in a manner strongly resembling our Fig. 5 (see Harding et al., 1999; Beniamini et al., 2020, 2023). While previous models of flare-induced spindown are phenomenologically similar to that proposed here, we emphasise that our predictions (i) arise self-consistently from magnetothermal outputs, (ii) do not rely on global events, and (iii) only work in CT configurations for typical birth periods. The second point is worth bearing in mind since, although extensive torque variability was observed in the lead-up and aftermath to the event, minimal changes were seen in $\dot{P}$ surrounding the giant flare from SGR 1806–20 (Younes et al., 2017). One way in which these scenarios could be confronted in future is through metal abundances: mass-loaded winds should invite $r$-process nucleosynthesis and hence enrich LPT nurseries (Patel et al., 2025).
@@
Another possibility proposed by Lander et al. (2026) for producing isolated LPTs is that of a delayed Meissner expulsion. This could resurrect a slow, CC magnetar by pushing flux into the crust. This scenario would, in essence, be represented by the left panel of Fig. 4 but with a lower ‘initial’ $E_{\rm rot}$, so that when crustal failures begin to sap energy the star could reach $\bar{\zeta}E^{\rm acc}_{\rm fail}>E_{\rm rot}$. However, forcing huge fluxes into the crust should generate heat: this fits with 1E 1613 but is difficult to reconcile with the coldest LPTs. One may also anticipate a skew towards larger burst energies in this case (top panel of Fig. 3), which is tension with the low energy of the DA J1832 event and lack of visible outbursts from other LPTs. If a bright flare from a LPT is detected in future, such a scenario may be favoured since the maximum failure energy of CT models is of order $\sim 10^{41}$ erg (bottom panel of Fig. 3). Hall waves of high amplitude may also be launched into the crust as flux is thrust there (Bransgrove et al., 2025), through a Meissner expulsion or otherwise. Notably, waves of the low-frequency, whistler variety may enduringly propagate in latitudinal directions such that the ambient poloidal field oscillates violently, instigating intermittent heating, anomalous torques, avalanche phenomena, and multiband transients. Ab initio modelling in this direction could pave the way for population synthesis studies of canonical vs. ‘late-blooming’ vs. ‘reborn’ magnetars.
@@ Acknowledgements
We thank Adam Dong, Zorawar Wadiasingh, Yuri Levin, and Andrei Beloborodov for valuable discussions, and Alex Cooper for insights on pulsar activity driven by plastic flow. AGS acknowledges funding from the European Union’s Horizon MSCA-2022 research and innovation programme “EinsteinWaves” under grant agreement No. 101131233, the Deutsche Forschungsgemeinschaft individual research grant 570901071, and the High Performance and Cloud Computing Group at the Zentrum für Datenverarbeitung of the University of Tübingen (Project “AnqaGW”). CD is supported by a Juan de la Cierva fellowship JDC2023‐052227‐I, funded by MINISTERIO DE CIENCIA, INNOVACIÓN Y UNIVERSIDADES, AGENCIA ESTATAL DE INVESTIGACIÓN, and UNIÓN EUROPEA (FSE+). We gratefully acknowledge the support provided by the Conselleria d’Educació, Cultura, Universitats i Ocupació de la Generalitat Valenciana through Prometeo Project CIPROM/2022/13, and from the AEI grant PID2021-127495NB-I00 funded by MCIN/AEI/10.13039/501100011033.
Author Contributions
The first two authors of this work (C. Dehman and A.G. Suvorov) contributed significantly to the conceptualisation, implementation, model design, writing, and editing: to that degree it is with regret that lead authorship cannot be more perceptibly shared. Their contributions were distinct but approximately equal (e.g., A.G. Suvorov handled spindown elements and C. Dehman conducted the magnetothermal simulations and formulated the corresponding results presented in this manuscript, with ongoing mutual feedback). J. Pons was also directly involved in matters of editing, presentation, and physics discussions regarding the key properties of the model(s) and simulation setup(s).
@@ Appendix A Numerical details on magnetothermal scheme
We incorporate superfluid and superconducting gaps for neutrons and protons, respectively, as they play a crucial role in cooling timescales by influencing the heat capacity and neutrino emissivity. While our model includes these effects, we do not alter the equations of motion to incorporate additional effects. For a comprehensive formulation of dissipative magnetohydrodynamic equations for finite-temperature superfluid and superconducting charged relativistic mixtures that account for particle diffusion, we refer to Dommes and Gusakov (2021). Thermal and electrical conductivities are computed using realistic data from the IOFFE repository [^3] http://www.ioffe.ru/astro/conduct/. For a complete discussion of all these microphysical ingredients, beyond that which we can adequately explore here, we refer the reader to Potekhin et al. (2015).
To describe the hydrostatics of the background state, we solve the Tolman-Oppenheimer-Volkov equations with the BSk24 equation of state (EOS), not accounting for (small) thermal corrections (Pearson et al., 2018). This EOS includes consistent crust and core regions and matches well with observational constraints from gravitational waves from binary-mergers, maximum-mass observations of pulsars, fast-cooling constraints, phase-coherent X-ray spectra mapped by the Neutron Star Interior Composition Explorer Mission (NICER), and other channels (Marino et al., 2024; Suvorov et al., 2024). At the magnetic field strengths and temperatures considered here, such a cold and unmagnetised EOS is an excellent approximation. A spherically-symmetric star is assumed, ignoring rotational oblateness and/or magnetic deformations to the mass-density. The above factors could, in principle, influence the evolution (at least at early times) and introduce complicated anisotropies into the crustal conductivity profiles, but such effects are expected to be small.